# Tripartite information of free fermions: a universal entanglement coefficient from the sine kernel

### 04 Mar 2026 v2

@@ Abstract
We study the tripartite information $I_{3}$ of free fermions on two-dimensional lattices partitioned into three adjacent strips of width $w$. Translation invariance yields the exact decomposition $I_{3}=\sum_{k_{y}}g(k_{F}(k_{y})\,w)$, where $g(z)$ is a universal function of the scaling variable $z=k_{F}w$, determined by the spectrum of the sine-kernel (Slepian) integral operator. We prove that $g(z)$ has a unique zero at $z^{*}=1.3288\pm 0.0001$: modes with $k_{F}w<z^{*}$ violate monogamy of mutual information ($g>0$), while modes with $k_{F}w>z^{*}$ satisfy it ($g<0$).
The central analytical result is $g(z)=cz+O(z^{3}\ln z)$ with $c=3\ln(4/3)/\pi\approx 0.2747$, derived from the rank-1 limit of the sine kernel. Two exact cancellations—of the $z\ln z$ area-law terms and of the $z^{2}$ terms—are intrinsic to the $I_{3}$ combination. The coefficient $c$ generalizes to $n$-partite information: $c_{n}=(n/\pi)\ln R_{n}$ with $R_{n}$ a rational number from binomial combinatorics. For Rényi entropy of index $\alpha$, we prove that $g_{\alpha}(z)\sim z^{\alpha}$ for $\alpha<2$ and $g_{2}(z)=-(8/\pi^{3})z^{3}$: von Neumann entropy ($\alpha=1$) uniquely gives linear sensitivity to Lifshitz transitions, while Rényi-2 gives only cubic sensitivity. We verify all predictions on square, triangular, and cubic lattices.
@@ 1 Introduction
(iii) A generalization to $n$-partite information ($c_{n}=(n/\pi)\ln R_{n}$) and to Rényi entropy, proving that von Neumann entropy is the unique index giving linear sensitivity to Lifshitz transitions.
@@ 2.3 Universal function $g(z)$
The key properties of $g(z)$, established by a combination of analytical arguments and computer-assisted verification (see Appendix A), are:
Proposition. (a) $g(0)=0$ and $g^{\prime}(0^{+})=c>0$. (b) $g$ has a unique maximum at $z_{\max}\approx 0.56$, with $g(z_{\max})\approx 0.088$. (c) $g$ has a unique zero at $z^{*}=1.3288\pm 0.0001$. (d) $g(z)<0$ for all $z>z^{*}$, with $g(z)\to(1/6)\ln(3/4)+O(e^{-cz})<0$ as $z\to\infty$.
The zero $z^{*}$ separates two regimes. For $z<z^{*}$, the direct mutual information I(A:D)I(A{:}D) across the gap $B$ exceeds the concavity deficit $\Delta^{2}S$ of single-interval entropy, giving $g>0$. For $z>z^{*}$, entropy concavity dominates, giving $g<0$. The zero is determined to $0.12\%$ by the cancellation of only two Slepian eigenvalue contributions, with a rigorous tail bound from prolate spheroidal eigenvalue estimates [10].
@@ 3.1 Rank-1 structure at small $z$
At small $z=k_{F}w$, the sine kernel becomes approximately constant: $\mathrm{sinc}(k_{F}d)=1+O(z^{2}/w^{2})$ for all site separations $d$ within the $3w$-site region. The correlation matrix reduces to rank 1, with a single nonzero eigenvalue $\lambda_{0}^{(n)}=nz/\pi$ for a block of $nw$ sites [10].
@@ 3.3 Higher-order structure
with $D\approx 0.218$, $E\approx-0.226$, and $\sum a_{k}k^{3}=3-24+27=6\neq 0$.
@@ 5.2 Numerical verification
Table 1: Rényi scaling of $g_{\alpha}(z)\sim z^{\beta}$. Exponents extracted from log-log fits at $z\in[10^{-3},10^{-2}]$ with $w=80$. The $\alpha=2$ coefficient $8/\pi^{3}\approx-0.258$ is confirmed to $0.15\%$.
@@ 5.4 Physical consequence
The pocket signal at a Lifshitz transition scales as $\Delta I_{3}\propto g^{\prime}(0^{+})\cdot\delta$. For von Neumann: $\Delta I_{3}^{(\alpha=1)}\propto c\,\delta$. For Rényi-2: $\Delta I_{3}^{(\alpha=2)}\propto\delta^{3}$three orders of magnitude weaker at small $\delta$. This means experimental probes measuring $S_{2}$ [16, 17] would miss the leading Lifshitz signal entirely.
@@ 6 Physical consequences
Neck transition (VHS). At a van Hove singularity, modes near the saddle have $k_{F}(q)=\gamma\sqrt{q^{2}+\delta/\beta}$, and $g^{\prime}(k_{F}w)\approx c$ for the small-$k_{F}$ modes:
@@
Entanglement tomography. The angular dependence $I_{3}(\theta)$ for strips at angle $\theta$ probes Fermi surface shape. For an $n$-fold deformation $k_{F}(\phi)=\bar{k}_{F}(1+\varepsilon\cos n\phi)$, the Widom coefficient responds at $O(\varepsilon^{2})$ (perimeter is quadratic in $\varepsilon$ at fixed area), but $I_{3}(\theta)$ responds at $O(\varepsilon)$ through the nonlinearity of $g(z)$. Thus $I_{3}$ detects Fermi surface deformations invisible to the leading entanglement entropy.
@@ 7.6 Pocket formula
7.7 Interactions: shift of $z^{*}$
Table 8: $z^{*}$ vs. interaction $V/t$ from exact diagonalization of the $t$-$V$ model ($L=24$, $w=2$).
$V/t$ $K$ $z^{*}$ $\Delta z^{*}/z^{*}_{0}$
$0.00$ $1.000$ $1.2184$
+1.00+1.00 $0.750$ $1.2109$ $0.62\%$
1.00-1.00 $1.500$ $1.1934$ $2.06\%$
The smallness of the shift ($\lesssim 2\%$ for $|V|\leq t$) reflects the near-discontinuity of $n(k)$ at $k_{F}$ in a Luttinger liquid with $K$ close to unity.
@@ 8 Discussion
Widom coefficient vs. $I_{3}$. The Widom coefficient $c_{1}$ in $S_{A}\sim c_{1}L\ln L$ [12, 13, 14] depends only on the FS perimeter. Two Fermi surfaces with the same perimeter but different curvatures are indistinguishable by $c_{1}$ but distinguishable by $I_{3}(\theta)$, which responds linearly to shape deformations. The $z^{2}$ cancellation ($\sum a_{k}k^{2}=0$) ensures that this sensitivity comes from the nonlinearity of $g(z)$ rather than trivial area scaling.
Open questions. Two problems emerge. First: the form of $g_{K}(z)$ for a Luttinger liquid with parameter $K$. The single-particle correlator decays as $\cos(k_{F}r)/r^{(K+K^{-1})/2}$, modifying the cross-gap mutual information that determines $z^{*}$. A systematic DMRG determination of $z^{*}(K)$ and $c_{K}$ on long chains is within reach. Second: the product $R_{n}$ in Eq. (10) may have a combinatorial interpretation—as a ratio of volumes or factorials—that connects the entanglement coefficient to classical enumeration.
@@ 9 Conclusions
We have established a complete analytical framework for the tripartite information of free fermions on two-dimensional lattices. The $k_{y}$-decomposition $I_{3}=\sum g(k_{F}(k_{y})\,w)$ reduces the problem to the universal function $g(z)$, whose unique zero at $z^{*}=1.329$ controls scale-dependent MMI. The small-$z$ coefficient $c=3\ln(4/3)/\pi$, derived from the rank-1 sine kernel, determines entanglement singularities at Lifshitz transitions. The $n$-partite generalization $c_{n}=(n/\pi)\ln R_{n}$ and the Rényi analysis—proving that von Neumann entropy uniquely gives linear Lifshitz sensitivity—complete the analytical picture.
Together, these results establish $I_{3}$ as a quantitative, direction-resolved probe of Fermi surface geometry with analytically controlled predictions.
@@ Appendix A Properties of $g(z)$: proof of the Proposition
(a) At $z=0$, all eigenvalues vanish, $h(0)=0$, so $g(0)=0$. For $z>0$: the rank-1 derivation (Section 3) gives $g(z)/z\to c$ as $z\to 0^{+}$.
(b) Verified by central differences at 200 points on $[0,0.60]$: $g^{\prime}(z)>0$ on $(0,0.56)$ and $g(0.56)\approx 0.088$.
(c) Verified $g^{\prime}(z)<0$ at 200 points on $[0.56,1.40]$. Combined with $g(0.56)>0$ and $g(1.40)<0$, the intermediate value theorem gives a unique zero $z^{*}\in(0.56,1.40)$.
(d) We establish $g(z)<0$ on $(z^{*},\infty)$ by Lipschitz-bound arguments on subintervals $z^{*},1.40]$, $[1.40,3.0]$, $[3.0,10.0]$ using 500, 200, and 250 grid points respectively. For $z>10$, the asymptotic $g(z)\to(1/6)\ln(3/4)<0$ follows from the Widom formula: linear terms cancel by $\sum a_{k}k(3-4-2+3)=0$ and logarithmic terms give $g_{\log}=(1/6)\ln(3/4)<0$.

### 08 Mar 2026 v3

@@ Abstract
We study the tripartite information $I_{3}$ of free fermions on two-dimensional lattices partitioned into three adjacent strips of width $w$. Translation invariance yields the exact decomposition $I_{3}=\sum_{k_{y}}g(k_{F}(k_{y})\,w)$, where $g(z)$ is a universal function of the scaling variable $z=k_{F}w$, determined by the spectrum of the sine-kernel (Slepian) integral operator. We find that $g(z)$ has a unique zero at $z^{*}=1.329\pm 0.001$: modes with $k_{F}w<z^{*}$ violate monogamy of mutual information ($g>0$), while modes with $k_{F}w>z^{*}$ satisfy it ($g<0$).
The central analytical result is $g(z)=cz+O(z^{3}\ln z)$ with $c=3\ln(4/3)/\pi\approx 0.2747$, derived from the rank-1 limit of the sine kernel. Two exact cancellations—of the $z\ln z$ area-law terms and of the $z^{2}$ terms—are intrinsic to the $I_{3}$ combination. The coefficient $c$ generalizes to $n$-partite information: $c_{n}=(n/\pi)\ln R_{n}$ with $R_{n}$ a rational number from binomial combinatorics. For Rényi entropy of index $\alpha$, we show that $g_{\alpha}(z)\sim z^{\alpha}$ for $\alpha<2$ and $g_{2}(z)=-(8/\pi^{3})z^{3}$: von Neumann entropy ($\alpha=1$) uniquely gives linear sensitivity to Lifshitz transitions, while Rényi-2 gives only cubic sensitivity. We verify all predictions on square, triangular, and cubic lattices. DMRG calculations on the interacting $t$-$V$ model provide evidence that the linear coefficient is independent of the Luttinger parameter $K$.
@@ 1 Introduction
(iii) A generalization to $n$-partite information ($c_{n}=(n/\pi)\ln R_{n}$) and to Rényi entropy, showing that von Neumann entropy is the unique index giving linear sensitivity to Lifshitz transitions.
@@ 2.3 Universal function $g(z)$
The key properties of $g(z)$, established by analytical arguments at small and large $z$ and by systematic numerical evaluation at intermediate $z$ (see Appendix A), are:
Result. (a) $g(0)=0$ and $g^{\prime}(0^{+})=c>0$ (analytical). (b) $g$ has a unique maximum at $z_{\max}\approx 0.56$, with $g(z_{\max})\approx 0.088$ (numerical). (c) $g$ has a unique zero at $z^{*}=1.329\pm 0.001$ (numerical; see below for error estimate). (d) $g(z)<0$ for all $z>z^{*}$ (numerical for $z\in(z^{*},10)$; analytical for $z\to\infty$).
The zero $z^{*}$ separates two regimes. For $z<z^{*}$, the direct mutual information I(A:D)I(A{:}D) across the gap $B$ exceeds the concavity deficit $\Delta^{2}S$ of single-interval entropy, giving $g>0$. For $z>z^{*}$, entropy concavity dominates, giving $g<0$. The zero is determined to $0.12\%$ by the cancellation of only two Slepian eigenvalue contributions, with exponentially decaying tail contributions bounded by prolate spheroidal eigenvalue estimates [10].
@@ 3.1 Rank-1 structure at small $z$
At small $z=k_{F}w$, the sine kernel becomes approximately constant: $\mathrm{sinc}(k_{F}d)=1+O(z^{2}/w^{2})$ for all site separations $d$ within the $3w$-site region. The correlation matrix reduces to rank 1, with a single nonzero eigenvalue $\lambda_{0}^{(n)}=nz/\pi$ for a block of $nw$ sites. This follows from the Slepian eigenvalue asymptotics $\lambda_{m}\sim(z/\pi)^{2m+1}$ for $z\to 0$ [10], which ensures exponential suppression of all $m\geq 1$ modes at small $z$.
@@ 3.3 Higher-order structure
with $D\approx 0.218$ and $E\approx-0.226$ (determined numerically), and $\sum a_{k}k^{3}=3-24+27=6\neq 0$.
@@ 5.2 Numerical verification
Table 1: Rényi scaling of $g_{\alpha}(z)\sim z^{\beta}$. Exponents extracted by linear fit of $\ln|g_{\alpha}(z)|$ vs. $\ln z$ over $z\in[10^{-3},10^{-2}]$ using Toeplitz matrices at $w=80$. Fit residuals are below $1\%$ for all $\alpha$. The $\alpha=2$ coefficient $8/\pi^{3}\approx-0.258$ is confirmed to $0.15\%$.
@@ 5.4 Physical consequence
The pocket signal at a Lifshitz transition scales as $\Delta I_{3}\propto g^{\prime}(0^{+})\cdot\delta$. For von Neumann: $\Delta I_{3}^{(\alpha=1)}\propto c\,\delta$. For Rényi-2: $\Delta I_{3}^{(\alpha=2)}\propto\delta^{3}$—suppressed by a factor $\delta^{2}$ relative to the von Neumann signal. This means experimental probes measuring $S_{2}$ [16, 17] would require correspondingly higher precision to detect the leading Lifshitz singularity.
@@ 6 Physical consequences
Neck transition (VHS). At a van Hove singularity, modes near the saddle have $k_{F}(q)=\gamma\sqrt{q^{2}+\delta/\beta}$, and for sufficiently small $w$ (so that all neck modes have $k_{F}w\ll z^{*}$), $g^{\prime}(k_{F}w)\approx c$:
@@
Entanglement tomography. The angular dependence $I_{3}(\theta)$ for strips at angle $\theta$ probes Fermi surface shape. For an $n$-fold deformation $k_{F}(\phi)=\bar{k}_{F}(1+\varepsilon\cos n\phi)$, the Widom coefficient responds at $O(\varepsilon^{2})$ (perimeter is quadratic in $\varepsilon$ at fixed area). In contrast, $I_{3}(\theta)$ is expected to respond at $O(\varepsilon)$ through the nonlinearity of $g(z)$, since the integral $\int g(k_{F}(\phi)\,w)\,d\phi$ is generically linear in $\varepsilon$ when $g$ is nonlinear. A quantitative analysis of this sensitivity is left for future work.
@@ 7 Numerical verification
All free-fermion results are computed via the Peschel formula applied to $3w\times 3w$ Toeplitz matrices at w=64w=648080. Finite-$w$ corrections are below $0.1\%$ for all quoted values, as verified by comparing $w=64$ and $w=128$. Lattice calculations use $L=256$ unless noted otherwise.
@@ 8 Discussion
Widom coefficient vs. $I_{3}$. The Widom coefficient $c_{1}$ in $S_{A}\sim c_{1}L\ln L$ [12, 13, 14] depends only on the FS perimeter and is always positive. In contrast, $I_{3}$ changes sign as the FS geometry varies: Table 3 shows that $I_{3}$ reverses from positive to negative as $t^{\prime}$ increases, detecting the disappearance of nested modes with $k_{F}w<z^{*}$ that the Widom coefficient cannot distinguish. More generally, $I_{3}(\theta)$ for strips at angle $\theta$ may be sensitive to FS shape deformations at $O(\varepsilon)$, while the Widom coefficient responds only at $O(\varepsilon^{2})$. The $z^{2}$ cancellation ($\sum a_{k}k^{2}=0$) ensures that this sensitivity comes from the higher-order structure of $g(z)$ rather than trivial area scaling.
Interactions. The rank-1 derivation of Section 3 relies on the correlation matrix becoming approximately constant over the block at small $z$. For free fermions, this follows from $\mathrm{sinc}(k_{F}d)\approx 1$ when $k_{F}d\ll 1$. In a Luttinger liquid, the single-particle correlator acquires an anomalous exponent, $\langle c^{\dagger}_{i}c_{j}\rangle\sim\sin(k_{F}r)/r^{(K+K^{-1})/2}$, and the rank-1 approximation no longer holds in general. However, the rank-1 eigenvalue $\lambda_{0}=\sum_{j}C_{0j}/w$ is determined by the integral of the correlator over the block, which equals the mean occupation $k_{F}/\pi$ regardless of the decay exponent; it is the two sum rules $\sum a_{k}k=\sum a_{k}k^{2}=0$ that then isolate this quantity as the sole contributor to $g(z)$ at leading order. DMRG calculations on the $t$-$V$ model at four system sizes (L=64L=64512512, w=2w=244, $K\in[0.75,1.50]$) show that the ratio $R(K)=I_{3}(\Delta)/I_{3}(0)$ converges toward unity at small $z$: |1R|=0.015|1-R|=0.015 at $z=0.39$, $L=512$. This is consistent with $c_{K}=c$ for all $K$, though a residual $K$-dependence at the percent level cannot be excluded from the present data. At larger $z$, the convergence is slower ($|1-R|\approx 0.07$ at $z=0.79$, $L=512$), and the full function $g_{K}(z)$ may acquire $K$-dependent corrections through the subleading Slepian eigenvalues. A definitive determination requires infinite-system DMRG or periodic boundary conditions to eliminate open-boundary artifacts.
Open question. The product $R_{n}$ in Eq. (10) may have a combinatorial interpretation—as a ratio of volumes or factorials—that connects the entanglement coefficient to classical enumeration.
@@ 9 Conclusions
We have established a systematic analytical framework for the tripartite information of free fermions on two-dimensional lattices. The $k_{y}$-decomposition $I_{3}=\sum g(k_{F}(k_{y})\,w)$ reduces the problem to the universal function $g(z)$, whose unique zero at $z^{*}=1.329$ controls scale-dependent MMI. The small-$z$ coefficient $c=3\ln(4/3)/\pi$, derived from the rank-1 sine kernel, determines entanglement singularities at Lifshitz transitions. The $n$-partite generalization $c_{n}=(n/\pi)\ln R_{n}$ and the Rényi analysis—showing that von Neumann entropy uniquely gives linear Lifshitz sensitivity—complete the analytical picture.
Together, these results establish $I_{3}$ as a quantitative, direction-resolved probe of Fermi surface geometry with analytically controlled predictions. DMRG calculations on the $t$-$V$ model provide evidence that $c$ is robust against interactions.
@@ Appendix A Properties of $g(z)$: proof of the Proposition
(a) At $z=0$, all eigenvalues vanish, $h(0)=0$, so $g(0)=0$. For $z>0$: the rank-1 derivation (Section 3) gives $g(z)/z\to c$ as $z\to 0^{+}$. This is analytical.
(b) Numerically verified by evaluating $g^{\prime}(z)$ via central differences at 200 equally spaced points on $[0,0.60]$ using Toeplitz matrices of size $3w$ with $w=64$: $g^{\prime}(z)>0$ on $(0,0.56)$ and $g(0.56)\approx 0.088$.
(c) Numerically verified that $g^{\prime}(z)<0$ at 200 points on $[0.56,1.40]$ with the same method. Combined with $g(0.56)>0$ and $g(1.40)<0$, the intermediate value theorem gives a unique zero $z^{*}\in(0.56,1.40)$. The zero is located by bisection: $z^{*}(w{=}32)=1.32829$, $z^{*}(w{=}64)=1.32868$, $z^{*}(w{=}128)=1.32878$. The quoted value $z^{*}=1.329\pm 0.001$ is a conservative estimate: the variation between $w=64$ and $w=128$ is $0.0001$, but the extrapolation to $w\to\infty$ introduces additional uncertainty of order $1/w\sim 0.01$, which we round conservatively.
(d) We verify $g(z)<0$ on $(z^{*},10)$ by evaluating $g$ on grids of 500, 200, and 250 points on $z^{*},1.40]$, $[1.40,3.0]$, $[3.0,10.0]$ respectively, with $w=64$. The maximum of $g$ on any subgrid is bounded by the grid value plus the grid spacing times a numerically estimated derivative bound. For $z>10$, the asymptotic $g(z)\to(1/6)\ln(3/4)<0$ follows analytically from the Widom formula: the linear terms cancel by $\sum a_{k}k=0$ and the surviving logarithmic terms give $g_{\log}=(1/6)\ln(3/4)<0$.